a steel rod near a blacksmith furnace is 1.00m long with a mass of 1.7kg and crosectional area 0.0001. if the rod absorb 40000j. find the increas in length of the rod.
1
Expert's answer
2020-10-30T13:19:27-0400

The equation of a termal balance here is:


"Q = mc\Delta T"

where "Q = 4\times 10^4J" is the amount of heat the rod has absorbed, "m = 1.7kg" is the mass of the rod, "c = 460J/(kg\cdot K)" is the specific heat capacity of steel and "\Delta T" is the increase in temperature. Thus, expressing the temperature increase, obtain:


"\Delta T = \dfrac{Q}{mc}"

According to the termal expansion law, the increas in length of the rod is given by:


"\Delta L = \alpha_L\cdot L\cdot \Delta T"

where "\alpha_L = 11\times 10^{-6}K^{-1}" is the coefficient of linear thermal expansion of steel, "L = 1m" is the initial length of the rod and "\Delta T" is the increase in temperature that we have already found. Substituting the expression for "\Delta T", obtain:


"\Delta L = \alpha_L\cdot L\cdot \dfrac{Q}{mc}\\
\Delta L = 11\times 10^{-6}\cdot 1\cdot \dfrac{4\times 10^4}{1.7\cdot 460} \approx 5.63\times 10^{-4}m"

Answer. "5.63\times 10^{-4}m".